### 15 Mar 2026 v1

# On Partition Functions for Time-Inhomogeneous Branching Random Walks

Abstract
We establish the phase transition and universality for the partition function of time-inhomogeneous branching random walks (BRWs) with decreasing variance increment, a model related to two-dimensional directed polymers. By modifying Berestycki’s GMC framework (removing martingale property dependency) and adapting it to discrete BRWs, we prove that the critical value of time-inhomogeneous BRWs coincides with that of time-homogeneous ones, and the partition functions converge in $L^{1}$ in the subcritical regime. We also extend the universality at the critical parameter, showing the same decay rate of partition functions. Our approach reveals the potential for such framework in GMC, which provides a new technical path for martingale-free processes and random fields beyond log-correlated. Finally, we raise some open problems related to GMC beyond log-correlations, branching Brownian motions and directed polymers.
1 Introduction
Branching random walks (BRWs) are a natural model in probability theory, motivated by their biological applications, that describe the evolution of a population of particles with spatial motion. Much is known about the phase transition of their partition function (the so-called Biggins martingale); see [17]. More recently, their time-inhomogeneous counterpart, first introduced in [14], initiates the study of the asymptotic behaviors of BRWs with increments that exhibit a macroscopic and monotonic change in variance, revealing differences in extreme value statistics. The series of works on this model raises a more fundamental question: How do environmental fluctuations impact the asymptotic behavior of BRWs? In the present paper, we attempt to answer this question from a qualitative perspective by studying its partition function.
The partition function of time-inhomogeneous BRWs with monotonically decreasing variance—our primary focus in the present paper—is motivated by and closely related to two-dimensional directed polymers (see Section 4 for a detailed explanation of this relationship). On the other hand, the model of Gaussian Multiplicative Chaos (GMC) has long been viewed as a continuous analogue of BRWs, and numerous works connect these two models: for instance, many results from one often inspire analogous results for the other, and many approaches used in BRWs can be applied to those in GMC-type objects (see, for example [18] for a survey). In the present paper, we aim to proceed in the opposite direction: we use recent developments in GMC to further the study of BRWs.
Nonetheless, the complexity of the underlying partition function is a more intricate topic, as the classical martingale structure no longer holds and the correlation function is not sufficiently regular. Such problems have been avoided in most literature on BRWs and directed polymers. However, recent developments in GMC extend the theory to the non-martingale construction of GMC measures. In particular, Berestycki’s paper [4] provides a novel perspective that focuses on analyzing the ”thick points” that dominate the total mass. Motivated by this framework, we can not only provide an alternate proof of the classical result regarding phase transition of BRWs but also perform a similar construction in the BRWs literature and establish the universal behavior of the partition functions: in the sub-critical regime, the two partition functions (for time-homogeneous and time-inhomogeneous environments) converge to the same non-degenerate random variable. Our technical framework not only unifies the phase transition analysis of homogeneous and time-inhomogeneous BRWs but also provides a reusable tool for martingale-free random processes, enriching the cross research between BRWs and GMC.
These counterintuitive results in the present paper suggests that, in such settings, the fundamental asymptotic property of the partition function is determined solely by the underlying branching process and proves robust to time inhomogeneity in the weight increments. The core of the approaches used in the present paper is the observation that the total weight is dominated by the ”early generations” of the BRW—a key universal property unreported for time-inhomogeneous BRWs, distinct from that in [14].
Looking ahead, the results in the present paper suggest that, within the framework of GMC, the correlation function of time-inhomogeneous BRWs falls outside the scope of log-correlation, which in turn implies that the GMC framework can be extended to more general settings. Additionally, motivated by the fundamental role of BRWs in various branching processes and directed polymer models, our study raises a broader question: How does time-inhomogeneity impact phase transitions in these systems? The authors believe that the framework discussed in the present paper can be extended to some of these models.
1.1 Outline:
In Section 2, we introduce the model of BRWs (in time-homogeneous and time-inhomogeneous environments), the GMC and their connections, and state our main results. In Section 3, we will prove Theorem 2.2, and we can see that Theorem 2.1 serves as a direct consequence of the proof. Then we will prove the upper bound of $\bar{\beta}_{c}$, and together it shows that $\bar{\beta}_{c}=\beta_{c}$. In Section 3.1, we extend the universality to the critical parameter, showing the same polynomial decay rate of partition functions for time-inhomogeneous BRWs 3.7. Finally, we will discuss our results in context and raise some potential directions for further study in Section 4.
1.2 Acknowledgment:
The author wishes to thank Prof. Clément Cosco for proposing this interesting model, for insightful discussions that led to Theorem 2 and Theorem 3.6, and for valuable comments on the early stages of this work. The author is also very grateful to Jiaqi Liu for discussions about this research in the context of branching processes.
2 The Model of Branching Random Walks and Main Results
In the present paper, we consider a branching random walk on the real line $\mathbb{R}$. Initially, a particle sits at the origin. Its children form the first generation; their displacements from the origin correspond to a point process on the line (which could also be described by a Galton-Watson process, abbreviated as G-W process or G-W trees). These children have children of their own (who form the second generation) and behave—relative to their respective positions—like independent copies of the initial particle, and so on. We will focus on time-homogeneous and time-inhomogeneous Gaussian Branching Random Walks (BRWs) on a supercritical Galton-Watson tree $\mathbb{T}$ rooted at $o\in\mathbb{T}$ with mean offspring distribution equal to $\mathbb{E}[D_{1}]=d>1$, and the detailed definition of the model will be introduced below. In the rest of the paper, we will use the survival probability of the G-W tree:
$\mathbf{P}^{*}=\mathbf{P}(\cdot|\text{non-extinction)}$
We assign to each vertex $\mathbf{x}\in\mathbb{T}$ an independent random variable $w(\mathbf{x})$, such that $\{w(\mathbf{x})\}_{\mathbf{x}\in\mathbb{T}}$ are i.i.d. $\mathcal{N}(0,1)$ (For $\sigma>0$, we denote $\mathcal{N}(\mu,\sigma^{2})$ to be the Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.). For each $\mathbf{x}\in\mathbb{T}$, we write $|\mathbf{x}|=n$ if $v$ belongs to the $n$-th generation of the Galton–Watson tree, and we let $(x_{1},\dots,x_{n})$ be the unique geodesic path from $o$ to $v$. Since the geodesic is unique we can refer to each branch at generation $n$ by the last vertex.
Let $W_{n}^{\beta}$ denote the (normalized) partition function for time-homogeneous BRWs at generation $n$, where $\beta\geq 0$ is a fixed parameter (which is so-called the inverse temperature in statistical mechanics), defined by
$W_{n}^{\beta}=\frac{1}{d^{n}}\sum_{|\mathbf{x}|=n}e^{\beta H_{n}(\mathbf{x})-\frac{1}{2}\beta^{2}n},$
where $H_{n}(\mathbf{x})=\sum_{i=1}^{n}w(x_{i})$ is the Hamiltonian associated with the branch $\mathbf{x}=(x_{1},\dots,x_{n})$. It is well known that the family of random variables $(W_{n}^{\beta})_{n\geq 0}$ is a unit-mean martingale, and thus converges almost surely to a nonnegative limit $W_{\infty}^{\beta}$. Applying Kolmogorov’s 0-1 law gives the usual dichotomy that exactly one of the events:
$\{W_{\infty}^{\beta}>0\}\text{ (weak disorder) and }\{W_{\infty}^{\beta}=0\}\text{ (strong disorder)}$
has probability 1 (a similar argument also applies to the partition function of time-inhomogeneous branching random walks, a model will be introduced later).
The following classical result of Biggins (see e.g. [17] for a more general statement and proof) characterizes the non-degeneracy of $W_{\infty}^{\beta}$:
Theorem 2.1.
(Phase Transition for Time-homogeneous BRWs) For a BRW on a supercritical Galton–Watson tree with mean offspring number $\mathbb{E}[D_{1}]=d>1$, there exists a phase transition for the non-degeneracy of $W_{\infty}^{\beta}$, i.e.,
$\beta_{c}:=\sup\Bigl\{\beta\geq 0\mid\mathbb{P}\{W_{\infty}^{\beta}=0\}>0\Bigr\}$
is well-defined and equals $\sqrt{2\log d}$.
The classical proof of the previous result (see, for example, Chapter 5 of [17]) relies on a so-called ”spine method” (or ”spine decomposition”), an extensively used technique in the branching process literature, where one uses a change-of-measure argument and tilts the measure towards a particle chosen uniformly at random. Leveraging the martingale property, it turns out that under the tilted measure, the chosen particle dominates the total mass.
Inspired by the long-standing relationship between BRWs and Gaussian Multiplicative Chaos (GMC), we provide a new proof of this result in the present paper by applying the analogous framework and approach from Berestycki’s paper [4] regarding the convergence of sub-critical GMC—this proof does not rely on the martingale property. The GMC framework in [4] is chosen due to its natural connection with BRWs (BRWs serve as the discrete analogue of continuous GMC) and its martingale-free property, which perfectly aligns with time-inhomogeneous BRWs (which lack a martingale structure) and will be introduced later in this section.
The relationship between BRWs and GMC can be briefly described as follows (roughly speaking, we embed a binary BRWs into the unit interval, which yields so-called ”multiplicative cascades”, a preliminary version of GMC): First, let $\tilde{I}$ denote the infinite union of sub-intervals of $[0,1]$ defined by
$\tilde{I}=\bigcup_{m\geq 1}\tilde{I}_{m}=\bigcup_{m\geq 1}\bigcup_{i=0}^{2^{m}-1}\left[\frac{i}{2^{m}},\frac{i+1}{2^{m}}\right],$
where $\tilde{I}_{m}$ denotes the set of $m$-th level sub-intervals. We then assign a random variable $W_{I}$ to each interval $I\in\tilde{I}$: for any $I\in\tilde{I}_{k}$ (with $k\in\mathbb{N}\cup\{0\}$), set $W_{I}=e^{\beta W_{I}^{\prime}-\frac{\beta^{2}}{2}}$ for some $\beta>0$, where $W_{I}^{\prime}\sim\mathcal{N}(0,1)$ and the $W_{I}^{\prime}$ are independent across distinct $I\in\tilde{I}$.
For any point $x\in[0,1]$, define $w_{n}(x)=\prod_{I\in\tilde{I},\,x\in I}W_{I}$ (equivalently, $w_{n}(x)=\prod_{I\in\tilde{I}}W_{I}\cdot\mathbf{1}_{\{x\in I\}}$, where $\mathbf{1}_{\{\cdot\}}$ denotes the indicator function). We then introduce the random measure
$\mu_{n}:=\mu(n,\beta)=w_{n}\cdot\lambda,$
where $\lambda$ is the Lebesgue measure on $[0,1]$. For each $n\geq 1$, $\mu_{n}$ defines a random measure (not necessarily a probability measure) on $[0,1]$, and by independence of $W_{I}$ we can see that $\mathbb{E}[\mu_{n}([0,1])]=\prod_{I\in\tilde{I}_{n}}\mathbb{E}[W_{I}]=1$ for all $n$. This construction can be extended to GW trees via analogous ideas.
The modern construction of GMC, originates from Kahane [13], can be roughly described as a random measure in a given domain $D\subset\mathbb{R}^{d}$, associated to a centered Gaussian field $h$ (a generalized function, not defined point-wise) and a constant parameter $\gamma>0$. Subsequently, such measures were generalized to a continuous setting on $\mathbb{R}^{d}$, yielding a Borel measure on a given domain $D\subset\mathbb{R}^{d}$ of the form
$\mu(dx)=e^{\gamma X(x)-\frac{1}{2}\gamma^{2}\mathbb{E}[X(x)^{2}]}\,\sigma(dx).$ (1)
In the expression above:
  • $h$ is a centered Gaussian generalized random field with covariance kernel $K(x,y)=-\log|x-y|+g(x,y)$, where $g(x,y)$ is continuous on $\bar{D}\times\bar{D}$;
  • $\gamma>0$ is a fixed constant;
  • $\sigma$ is a Radon measure on $\mathbb{R}^{d}$, serving as the reference measure.
In Berestycki’s paper [4] provides an elementary proof for the convergence of approximating measures $\mu_{\epsilon}$ constructed via the convolution of $h(x)$ with certain regularizing kernels $\theta_{\epsilon}$ to a non-trivial limiting measure $\mu$ in the sub-critical regime (specifically, when the parameter satisfies $\gamma<\sqrt{2d}$) and further demonstrates that this limiting measure $\mu$ is independent of the choice of regularizing kernels. The GMC framework with thick point analysis, i.e. the points $x\in D$ satisfying:
$\lim_{\epsilon\to 0}\frac{X_{\epsilon}(x)}{|\log 1/\epsilon|}=\gamma$
which was shown in [4] that they dominates the total GMC measure, and measure tilting is a more suitable technical tool for phase transition analysis for BRWs in time-inhomogeneous environments. In time-inhomogeneous environment, the weight increment of any branch $\mathbf{|x|}=n$ exhibits a macroscopic variation: let $f:[0,1]\to[0,1]$ be continuous at 0, monotone decreasing function satisfying $f(0)=1$ and $f(1)=0$. For each generation in G-W tree $h\in\{0,...,n\}$, the weight increment at step $h$ follows the distribution of $\sqrt{f(\frac{h}{n})}w(x_{h})$. In other words, the Hamiltonian of a branch $\mathbf{x}=\{x_{1},...,x_{n}\}$, denoted by $\bar{H}_{n}(\mathbf{x})$, is given by: $\bar{H}_{n}(x)=\sum_{i=1}^{n}\sqrt{f(\frac{i}{n})}w(x_{i})$, so that $\bar{H}_{n}(\mathbf{x})\sim\mathcal{N}(0,\sum_{i=1}^{n}f(\frac{i}{n}))$.
Analogous to the normalized partition function for the BRW with homogeneous weight increments, we define the normalized partition function for the time-inhomogeneous BRW of length $n$, denote $\bar{W}_{n}^{\beta}$, as follows:
$\bar{W}_{n}^{\beta}=\frac{1}{d^{n}}\sum_{\mathbf{|x|}=n}e^{\beta\bar{H}_{n}(x)-\frac{\beta^{2}}{2}\sum_{i=1}^{n}f(\frac{i}{n})w(x_{i})}$
However, unlike $(W_{n}^{\beta})_{n\geq 0}$ in time-homogeneous environment, the partition function for time-inhomogeneous environment is not a martingale (due to the non-constancy of $f$), so the convergence of $(W_{n}^{\beta})_{n\geq 0}$ is not automatic.
In the present paper, we will establish the following stronger result: when $\beta<\beta_{c}$, the difference $\bar{W}_{n}^{\beta}-W_{n}^{\beta}\to 0$ in the $L^{1}$-norm, which in turn implies convergence to 0 in probability. Notably, $\bar{W}_{\infty}^{\beta}$ follows the same law as $W_{\infty}^{\beta}$-the a.s. limit of partition function in homogeneous weights. As a direct consequence, we can see that $\bar{W}_{\infty}^{\beta}$ is almost surely non-negative when $\beta<\sqrt{2\log d}$:
Theorem 2.2.
(Phase Transition and Universality for Time-inhomogeneous BRW) Let $f:[0,1]\to[0,1]$ be a continuous, non-increasing function with $f(0)=1$. Define the normalized partition function for time-inhomogeneous weights by
$\bar{W}_{n}^{\beta}=\frac{1}{d^{n}}\sum_{|\mathbf{x}|=n}e^{\beta\bar{H}_{n}(\mathbf{x})-\frac{\beta^{2}}{2}\sum_{i=1}^{n}f\!\left(\frac{i}{n}\right)},$
where
$\bar{H}_{n}(\mathbf{x})=\sum_{i=1}^{n}\sqrt{f\!\left(\frac{i}{n}\right)}w(x_{i})$
is the Hamiltonian for a branch $\mathbf{x}$ with $|\mathbf{x}|=n$. Then $\bar{W}_{n}^{\beta}$ exhibits a phase transition for its limit in probability (whose existence is established separately), with critical value
$\bar{\beta}_{c}:=\sup\Bigl\{\beta\geq 0\mid\mathbb{P}\{\bar{W}_{\infty}^{\beta}=0\}>0\Bigr\},$
which is well-defined and equals $\sqrt{2\log d}$ (coincides with that of the homogeneous model $W_{n}^{\beta}$, i.e., $\beta_{c}=\bar{\beta}_{c}$). Moreover, the following universality in $L^{1}$ limit holds:
$\mathbb{E}\bigl[|W_{n}^{\beta}-\bar{W}_{n}^{\beta}|\bigr]\to 0\quad\text{as }n\to\infty,\text{ for all }\beta<\beta_{c}=\sqrt{2\log d}.$
Remark: Indeed, from the proof in Section 3 we can see that the restriction of $f$ in Theorem 2.2 could be relaxed: the proof will hold as long as $f:[0,1]\to[0,1]$ with $f(0)=1$ and being continuous at 0 (so it serves as a generalization of the condition of the decreasing variance in [14]), and the universality could be easily extended to any finite dimensional marginals.
Finally, the exact asymptotic behavior at criticality was established in the influential work of Hu and Shi,who proved that at the critical parameter $\beta_{c}$, $W_{n}^{\beta}$ possess the asymptotic behavior:
$W_{n}^{\beta}=n^{-\frac{1}{2}+o(1)}\text{ a.s. }$
Our next theorem shows that the universality principle extends to the critical parameter. This result was also motivated by the paper [2],which studies the decay rate of partition function near criticality:
Theorem 2.3.
(Asymptotics of partition function at criticality) At the critical parameter $\beta_{c}$, we have $\bar{W}_{n}^{\beta_{c}}=n^{-\frac{1}{2}+o(1)}$ almost surely.
The polynomial decay rate for time-inhomogeneous BRWs at criticality originates from the universal domination of early generations, which is consistent with time-homogeneous BRWs and reflects the insensitivity of phase transition to variance inhomogeneity.
3 Proof of Main Results
In this section, we establish the existence of critical parameters governing the phase transitions of the limit $\bar{W}_{\infty}^{\beta}$, denoted respectively $\bar{\beta_{c}}$, and further derive its exact value.
Before we dive into the proof, we would like to illustrate a standard summation argument (the $L^{2}$-computation) used in our analysis, which usually forms the main technicality in various research in branching processes, and will be used repeatedly (in various forms) throughout the paper. For simplicity, we would first focus on our discussion in the time-homogeneous weights, and it would be straightforward to extend the computation to the time-inhomogeneous weights. The key idea is to fix a branch $\mathbf{|x|}=n$ and sum over all possible depths for $|\mathbf{y}|=n$ at which the pair of branches $\mathbf{x}$ and $\mathbf{y}$ diverge.
Within the BRWs framework, the covariance $\text{Cov}\left(H_{n}(\mathbf{x}),H_{n}(\mathbf{y})\right)$ for two branches $|\mathbf{x}|=|\mathbf{y}|=n$ depends exclusively on their branching depth $h\in\{0,\dots,n\}$—the depth at which they first diverge. Formally, for branches $\mathbf{x}=(x_{1},\dots,x_{n})$ and $\mathbf{y}=(y_{1},\dots,y_{n})$, define their coincidence depth as:
$C(\mathbf{x},\mathbf{y}):=\sup\left\{k\geq 0\mid x_{k}=y_{k}\right\}$
with the convention that $C(\mathbf{x},\mathbf{y})=0$ if the set is empty. For such pair of branches, the covariance of their Hamiltonian satisfies $\text{Cov}\left(H(\mathbf{x}),H(\mathbf{y})\right)=2h$.
We fix a branch $\mathbf{|x|}=n$ (noting there are expected $d^{n}$ such branches) and sum over all possible branching depths $h\in\{0,\dots,n\}$ relative to $\mathbf{x}$. For each fixed $\mathbf{x}$ and depth $h$, the number of branches $\mathbf{y}$ satisfying $C(\mathbf{x},\mathbf{y})=h$ has expected value $d^{n-h}$, corresponding to the number of branches in a sub-Galton-Watson tree with (nh)(n-h) generations.
This structure allows us to rewrite the $L^{2}$-norm of $W_{n}^{\beta}$ as:
$\mathbb{E}\left[(W_{n}^{\beta})^{2}\right]=d^{-2n}\mathbb{E}\left[\sum_{\mathbf{|x|=n}}\sum_{h=0}^{n}\sum_{C(\mathbf{x,\mathbf{y}})=h}[e^{-n\beta^{2}+\beta(H_{n}(\mathbf{x})+H_{n}(\mathbf{y}))}]\right]$
by independence of the law of GW tree and the weight functions, we have:
$\mathbb{E}\left[(W_{n}^{\beta})^{2}\right]=\sum_{h=0}^{n}d^{-h}\mathbb{E}[e^{-n\beta^{2}+\beta(H(\mathbf{x})+H(\mathbf{y}))}]$
Since the Laplace transform of $H_{n}(\mathbf{x})+H_{n}(\mathbf{y})$ for the sum of Hamiltonian for two branches $\mathbf{x},\mathbf{y}$ where $C(\mathbf{x},\mathbf{y})=h$ is given by:
$\mathbb{E}\left[e^{\beta(H_{n}(\mathbf{x})+H_{n}(\mathbf{y}))}\right]=\mathbb{E}\left[e^{2\beta(\tilde{H}(\mathbf{x}))}\right]\mathbb{E}\left[e^{\beta(H^{\prime}(\mathbf{x}))}\right]\mathbb{E}\left[e^{\beta(H^{\prime}(\mathbf{y}))}\right]$
where $\tilde{H}(\mathbf{x})$ has law $\mathcal{N}(0,h)$ and $H^{\prime}(\mathbf{x}),H^{\prime}(\mathbf{y})$ both have law $\mathcal{N}(0,(n-h))$ and they are all independent, so this equals to:
$\mathbb{E}\left[e^{\beta(H(\mathbf{x})+H(\mathbf{y}))}\right]=e^{2\beta^{2}h}(e^{\frac{\beta^{2}}{2}(n-h)})^{2}=e^{\beta^{2}(n+h)}$
As a result, the previous $L^{2}$ norm of $W_{n}^{\beta}$ equals to:
$\mathbb{E}\left[(W_{n}^{\beta})^{2}\right]=\sum_{h=0}^{n}d^{-h}e^{-n\beta^{2}}e^{\beta^{2}(n+h)}=\sum_{h=0}^{n}e^{h(\beta^{2}-\log d)}$
and finally we can see that the series converges if and only if $\beta^{2}-\log d<0$, i.e., $\beta<\sqrt{\log d}$.
Then we will establish the existence of the $L^{2}$ critical parameter for the time-inhomogeneous weights, denoted $\bar{\beta}_{2}$:
Theorem 3.1.
(Time-inhomogeneous $L^{2}$ Critical Parameter $\bar{\beta}_{2}$)
The $L^{2}$-critical parameter in time-inhomogeneous environment, defined by:
$\bar{\beta}_{2}:=\sup_{\beta\geq 0}\left\{\sup_{n\geq 0}\mathbb{E}\left[\left(W_{n}^{\beta}\right)^{2}\right]<\infty\right\}$
exists and equals to $\bar{\beta}_{2}=\sqrt{\log d}$ (in particular, the $L^{2}$ critical parameters in both cases coincide).
Proof.
Following the previous computations, we may express the definition of $\bar{\beta}_{2}$ in terms of the $L^{2}$-norm of $\bar{W}_{n}^{\beta}$ as follows:
$\mathbb{E}\left[\left(\bar{W}_{n}^{\beta}\right)^{2}\right]=\sum_{h=0}^{n}d^{-h}e^{\beta^{2}\sum_{i=1}^{h}f\left(\frac{i}{n}\right)}$ (2)
We can easily see from the expression above that the right hand side is monotonic in $\beta$, hence well-defined. In both cases in the assertion, we will prove the upper bound and lower bound of $\bar{\beta}_{2}$ respectively.
For the lower bound: From the expression, we can see that each term in the series expression (2) for $\mathbb{E}\left[\left(\bar{W}_{n}^{\beta}\right)^{2}\right]$ is bounded above by the corresponding term in the time-homogeneous case (i.e., the terms in the series expansion $\mathbb{E}\left[\left(W_{n}^{\beta}\right)^{2}\right]$). As a result, we have when $\beta<\beta_{2}$, $\mathbb{E}\left[\left(\bar{W}_{n}^{\beta}\right)^{2}\right]$ is uniformly bounded in $n$.
For the upper bound: Fix $0<\epsilon<1$ and take $\beta>\sqrt{\log d}$. By the continuity of $f$, there exists $\delta>0$ such that $f(x)<1-\delta$ for all $0\leq x<\epsilon$. We then have:
$\mathbb{E}\left[\left(\bar{W}_{n}^{\beta}\right)^{2}\right]\geq\sum_{h=0}^{\lfloor\epsilon n\rfloor}d^{-h}e^{\beta^{2}\sum_{i=1}^{h}f\left(\frac{i}{n}\right)}$
$\geq\sum_{h=0}^{\lfloor\epsilon n\rfloor}Cd^{-h}e^{\beta^{2}h(1-\delta)}=\sum_{h=0}^{\lfloor\epsilon n\rfloor}Ce^{h\left(-\log d+\beta^{2}(1-\delta)\right)}$
where $C$ is a constant depending only on $\beta$. For any fixed $\beta>\sqrt{\log d}$, we can choose $\delta>0$ sufficiently small such that $\log d+\beta^{2}(1-\delta)>0$. By the continuity of $f$, $\delta$ can be made arbitrarily small by taking $\epsilon$ small enough, and the series diverges as $n\to\infty$. ∎
Prior to proceeding with the proofs, we state a key lemma that facilitates the transformation of Gaussian random variables across different measures. This result will play a central role in our proof (we refer readers to [3] , Lemma 2.5 for a proof):
Lemma 3.2.
(Cameron-Martin-Girsanov formula) Let $X=(X_{1},...,X_{n})$ be a Gaussian vector under the law $\mathbb{P}$, with mean $\mu$ and covariance matrix $V$. Let $\alpha\in\mathbb{R}^{n}$ be a vector and we define a new probability measure by:
$\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}=\frac{e^{\langle\alpha,X\rangle}}{\mathbb{E}[e^{\langle\alpha,X\rangle}]}$
(here $\langle\alpha,X\rangle$ is the scalar product of two vectors $\alpha$ and $X$) Then under $\tilde{\mathbb{P}}$, $X$ is still a Gaussian vector, with covariance matrix $V$ and mean $\mu+V\alpha$.
In other (more plain) words, suppose we tilt the law of a Gaussian vector $X=(X_{1},...,X_{n})$ by some linear functional $X\to\langle\alpha,X\rangle$. Then the process remains Gaussian, with unchanged covariances, however the mean is shifted, and the new mean of the random variable $X_{i}$ (which original law under $\mathbb{P}$ is $X_{i}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$, is:
$\mu_{i}^{\prime}=\mu_{i}+\text{Cov}(X_{i},\langle\alpha,X\rangle)$
Having the lemma stated, we proceed our proof of the $L^{1}$ convergence in the sub-critical regime.
Theorem 3.3.
(Phase Transition and Universality for Time-inhomogeneous BRW) Let $f:[0,1]\to[0,1]$ be a continuous, non-increasing function with $f(0)=1$. Define the normalized partition function for time-inhomogeneous weights by
$\bar{W}_{n}^{\beta}=\frac{1}{d^{n}}\sum_{|\mathbf{x}|=n}e^{\beta\bar{H}_{n}(\mathbf{x})-\frac{\beta^{2}}{2}\sum_{i=1}^{n}f\!\left(\frac{i}{n}\right)},$
where
$\bar{H}_{n}(\mathbf{x})=\sum_{i=1}^{n}\sqrt{f\!\left(\frac{i}{n}\right)}w(x_{i})$
is the Hamiltonian for a branch $\mathbf{x}$ with $|\mathbf{x}|=n$. Then the following universality in $L^{1}$ limit holds:
$\mathbb{E}\bigl[|W_{n}^{\beta}-\bar{W}_{n}^{\beta}|\bigr]\to 0\quad\text{as }n\to\infty,\text{ for all }\beta<\beta_{c}=\sqrt{2\log d}.$
As a direct consequence, when $\beta<\beta_{c}$, $\bar{W}_{n}^{\beta}$ converges to an a.s. non-degenerate limit $\bar{W}_{\infty}^{\beta}$ in probability, which has the same law as $W_{\infty}^{\beta}$.
Proof.
We prove the $L^{1}$-convergence by dividing the branches into ”good environment” and its complement, for both time-homogeneous and time-inhomogeneous environments. Using the Cameron-Martin-Girsanov formula mentioned previously, we can see that the branches in ”good environment” in both cases dominates the total mass (i.e. the $L^{1}$-norm), and converges in $L^{2}$ to the same random variable.
Let $\beta<\alpha<2\beta$ be a fixed constant. For any branch $|\mathbf{x}|=n$, we define the ”good environment” event for the time-homogeneous BRWs, denoted $G_{n}^{\alpha}(\mathbf{x})$ (roughly speaking, it corresponds to the environment where the Hamiltonian of a branch $\mathbf{x}$ remains ”controlled”) for all $T\in\{n_{0},\ldots,n\}$ ,with $n_{0}\geq 0$ a fixed integer, by:
$G_{n}^{\alpha}(\mathbf{x}):=\left\{\forall T\in\{n_{0},\ldots,n\},\,H_{T}(\mathbf{x})<\alpha T\right\}.$
We define $J_{n}$ as the partition function restricted to good environments (since there is no ambiguity, we drop $\beta$ in the notation):
$J_{n}=\frac{1}{d^{n}}\sum_{|\mathbf{x}|=n}e^{\beta H_{n}(\mathbf{x})-\frac{1}{2}\beta^{2}n}\mathbf{1}_{\{\mathbf{x}\in G_{n}^{\alpha}(\mathbf{x})\}},$
and $K_{n}$ as its complement relative to $W_{n}^{\beta}$:
$K_{n}=\frac{1}{d^{n}}\sum_{|\mathbf{x}|=n}e^{\beta H_{n}(\mathbf{x})-\frac{1}{2}\beta^{2}n}\mathbf{1}_{\{\mathbf{x}\in(G_{n}^{\alpha}(\mathbf{x}))^{c}\}}.$
This gives the $L^{1}$-norm decomposition:
$\mathbb{E}\left[W_{n}^{\beta}\right]=\mathbb{E}\left[J_{n}\right]+\mathbb{E}\left[K_{n}\right].$
Analogously, we denote by $\bar{G}_{n}$ $\bar{J}_{n}$ and $\bar{K}_{n}$ the above quantities for time-inhomogeneous BRWs, which are given by:
$\bar{G}_{n}^{\alpha}(\mathbf{x})=\{\forall T\in\{n_{0},...,n\},\bar{H}_{T}(\mathbf{x})<\alpha\sum_{i=1}^{T}f(\frac{i}{n})\}$
and
$\bar{J}_{n}=\frac{1}{d^{n}}\sum_{|\mathbf{x}|=n}e^{\beta\bar{H}_{n}(\mathbf{x})-\frac{\beta^{2}}{2}\sum_{i=1}^{n}f\left(\frac{i}{n}\right)w(x_{i})}\mathbf{1}_{\{\mathbf{x}\in\bar{G}_{n}^{\alpha}(\mathbf{x})\}}$
By the triangle inequality, we have:
$\mathbb{E}\left[|\bar{W}_{n}^{\beta}-W_{n}^{\beta}|\right]=\mathbb{E}\left[|J_{n}-\bar{J}_{n}+K_{n}-\bar{K}_{n}|\right]\leq\mathbb{E}\left[|J_{n}-\bar{J}_{n}|\right]+\mathbb{E}\left[K_{n}\right]+\mathbb{E}\left[\bar{K}_{n}\right].$
It follows that when $n_{0}\to\infty$, both $\mathbb{E}[K_{n}]$ and $\mathbb{E}[\bar{K}_{n}]$ tend to 0 uniformly in $n$. To verify this, define the tilted probability measure $\hat{\mathbb{P}}$ by:
$\frac{\text{d}\hat{\mathbb{P}}}{\text{d}\mathbb{P}}=\frac{e^{\beta\bar{H}_{n}(\mathbf{x})}}{\mathbb{E}\left[e^{\beta\bar{H}_{n}(\mathbf{x})}\right]}.$
[…]
$\hat{\mathbb{P}}\{\bar{G}_{n}^{\alpha}(\mathbf{x})\}=\hat{\mathbb{P}}\{\forall T\in\{n_{0},...,n\},\bar{H}_{T}(\mathbf{x})<\alpha\sum_{i=1}^{T}f\left(\frac{i}{n}\right)\}$
$\mathbb{P}\{\forall T\in\{n_{0},...,n\},\bar{H}_{T}(\mathbf{x})<(\alpha-\beta)\sum_{i=1}^{T}f\left(\frac{i}{n}\right)\}\geq\mathbb{P}\{\bar{G}_{n}^{\alpha-\beta}(\mathbf{x})\}$
Consequently,
[…]
where $\epsilon(n_{0})\to 0$ as $n_{0}\to\infty$ uniformly in $n$. The probability that any single branch $\mathbf{x}$ with $|\mathbf{x}|=n$ does not lie in the good environment can be made arbitrarily small by choosing $n_{0}$ sufficiently large, with the upper bound uniform in $n$:
$\mathbb{P}\left\{G_{n}^{\alpha}(\mathbf{x})\right\}\geq 1-\epsilon(n_{0}),$
where $\epsilon$ depends only on $\alpha$ and $n_{0}$, and satisfies $\epsilon(n_{0})\to 0$ as $n_{0}\to\infty$ for each fixed $\alpha>\beta$. To verify this, we bound the probability of the complement event using Gaussian tail estimates:
$\mathbb{P}\left\{\left(G_{n}^{\alpha}(\mathbf{x})\right)^{c}\right\}\leq\mathbb{P}\left\{\exists T\geq n_{0},\,H_{T}(\mathbf{x})\geq\alpha\sum_{i=1}^{T}f\left(\frac{i}{n}\right)\right\}\leq\sum_{T\geq n_{0}}e^{-\alpha^{2}\sum_{i=1}^{T}f\left(\frac{i}{n}\right)},$ (3)
which tends to 0 as $n_{0}\to\infty$ uniformly in $n$. A similar argument yields $\mathbb{E}[\bar{J}_{n}]=1-\epsilon^{\prime}(n_{0})$ where $\epsilon^{\prime}(n_{0})\to 0$ as $n_{0}\to\infty$ uniformly in $n$. While the right-hand side of (3) converges for any $\alpha>0$, we later use the good environment $G_{n}^{\alpha-\beta}(\mathbf{x})$ with $\alpha<2\beta$, justifying our restriction to $\beta<\alpha<2\beta$.
Next, we will show that when $\beta<\sqrt{2}\bar{\beta_{2}}$, $J_{n}$ and $\bar{J}_{n}$ are both uniformly bounded in $L^{2}$. Since $\bar{\beta}_{2}=\beta_{2}$, we have that $\bar{J}_{n}$ is uniformly bounded in $L^{2}$ when $\beta<\beta_{c}$. For pairs $(\mathbf{x},\mathbf{y})$ with $C(\mathbf{x},\mathbf{y})=h$, define the transformed measure $\tilde{\mathbb{P}}^{h}$ by:
$\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}=\frac{e^{\beta\left(\bar{H}_{n}(\mathbf{x})+\tilde{\bar{H}}_{n}(\mathbf{x})\right)}}{\mathbb{E}\left[e^{\beta\left(\bar{H}_{n}(\mathbf{x})+\tilde{\bar{H}}_{n}(\mathbf{x})\right)}\right]},$
and under $\tilde{\mathbb{P}}^{h}$, $\bar{H}_{h}(\mathbf{x})\sim\mathcal{N}(2\beta\sum_{i=1}^{h}f(\frac{i}{n}),\beta\sum_{i=1}^{h}f(\frac{i}{n}))$, so we have:
$\mathbb{E}\left[J_{n}^{2}\right]=\sum_{h=0}^{n}d^{-h}e^{\beta^{2}\sum_{i=1}^{h}f(\frac{i}{n})}\tilde{\mathbb{P}}^{h}\left\{\bar{G}_{n}^{\alpha}(\mathbf{x})\cap\tilde{\bar{G}}_{n}^{\alpha}(\mathbf{x})\right\}.$ (4)
$\leq\sum_{h=0}^{n}d^{-h}e^{\beta^{2}\sum_{i=1}^{h}f(\frac{i}{n})}\tilde{\mathbb{P}}^{h}\{\bar{H}_{h}(\mathbf{x})<\alpha\sum_{i=1}^{h}f(\frac{i}{n})\}\leq\sum_{h=0}^{n}d^{-h}e^{(\beta^{2}-\frac{1}{2}(\alpha-2\beta)^{2})\sum_{i=1}^{h}f(\frac{i}{n})}$ (5)
letting $\alpha\to\beta$, we can see that the series is uniformly bounded in $n$ when $\beta<\sqrt{2}\bar{\beta_{2}}$, by a direct comparison with the $L^{2}$ computation of $\bar{W}_{n}^{\beta}$ in equation 2. A similar comparison with the $L^{2}$ computation of $W_{n}^{\beta}$ yields that when $\beta<\sqrt{2}\beta_{2}=\beta_{c}$, $J_{n}$ is uniformly bounded in $L^{2}$.
It remains to show $J_{n}-\bar{J}_{n}\to 0$ in $L^{1}$, which we establish by proving that the convergence holds in $L^{2}$. We start by expressing the $L^{2}$-norm of $J_{n}-\bar{J}_{n}$ as
$\mathbb{E}\left[(\bar{J}_{n}^{\beta}-J_{n}^{\beta})^{2}\right]=\mathbb{E}\left[(\bar{J}_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right]+\mathbb{E}\left[(J_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right].$ (6)
We will show that for any fixed $n_{0}>0$, both terms on the right-hand side of (6) tend to 0 as $n\to\infty$.
Fix $n_{1}>0$; we decompose $\mathbb{E}\left[(\bar{J}_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right]$ by $C(\mathbf{x},\mathbf{y})=h\in\{0,...,n\}$ into two sums over $h\leq n_{1}$ and $h>n_{1}$, so that $\mathbb{E}\left[(\bar{J}_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right]$ equals:
$\sum_{h=0}^{n_{1}}d^{-h}[\left(e^{\beta^{2}\sum_{i=1}^{h}f\left(\frac{i}{n}\right)}\mathbb{P}{\{\bar{G}_{n}^{\alpha}(\mathbf{x})\cap\bar{\tilde{G}}_{n}^{\alpha}(\mathbf{x})\}}\right)$
$\left(e^{\frac{1}{2}\beta^{2}\left(\sum_{i=1}^{h}\left(1+\sqrt{f\left(\frac{i}{n}\right)}\right)^{2}-(h)-\sum_{i=1}^{h}f\left(\frac{i}{n}\right)\right)}\mathbb{P}{\{\bar{G}_{n}^{\alpha}(\mathbf{x})\cap\bar{\tilde{G}}_{n}^{\alpha}(\mathbf{x})\}}\right)]$
$\sum_{h=n_{1}+1}^{n}[(d^{-h}\left(e^{\beta^{2}\sum_{i=1}^{h}f\left(\frac{i}{n}\right)}\mathbb{P}{\{\bar{G}_{n}^{\alpha}(\mathbf{x})\cap\bar{\tilde{G}}_{n}^{\alpha}(\mathbf{x})\}}\right)$
$\left(e^{\frac{1}{2}\beta^{2}\left(\sum_{i=1}^{h}\left(1+\sqrt{f\left(\frac{i}{n}\right)}\right)^{2}-(h)-\sum_{i=1}^{h}f\left(\frac{i}{n}\right)\right)}\mathbb{P}{\{\bar{G}_{n}^{\alpha}(\mathbf{x})\cap\bar{\tilde{G}}_{n}^{\alpha}(\mathbf{x})\}}\right)]$
For $\beta<\beta_{c}$ and any fixed $n_{1}>0$, form Lemma 3.4 we can see that $J_{n}$ and $\bar{J}_{n}$ converges to the same random variable $L^{2}$, since $\mathbb{E}[(\bar{J}_{n})^{2}]$ and $\mathbb{E}[\bar{J}_{n}J_{n}]$ has the following expansions:
$\mathbb{E}\left[(\bar{J}_{n})^{2}\right]=\sum_{h=0}^{n_{1}}d^{-h}e^{\beta^{2}\sum_{i=1}^{h}f(\frac{i}{n})}g^{\alpha}(h)+F(n_{1})$
and
$\quad\mathbb{E}\left[\bar{J}_{n}J_{n}\right]=\sum_{h=0}^{n_{1}}d^{-h}e^{\beta^{2}\left(\sum_{i=1}^{h}\left(1+\sqrt{f\left(\frac{i}{n}\right)}\right)^{2}-(h)-\sum_{i=1}^{h}f\left(\frac{i}{n}\right)\right)}g^{\alpha}(h)+F^{\prime}(n_{1})$
where $F(n_{1})$ and $F^{\prime}(n_{1})$ can be arbitrarily small by choosing $n_{1}$ sufficiently large. Since $n_{1}>0$ is arbitrary, both $\mathbb{E}[(\bar{J}_{n})^{2}]$ and $\mathbb{E}[\bar{J}_{n}J_{n}]$ converge to $\sum_{h=0}^{\infty}d^{-h}e^{\beta^{2}h}g^{\alpha}(h)$ as $n\to\infty$.
To confirm this, note that for any fixed $n_{1}>0$ and $h\in\{0,1,\dots,n_{1}\}$, $\frac{h}{n}\to 0$ as $n\to\infty$; by continuity of $f$, this implies $f\left(\frac{h}{n}\right)\to f(0)=1$ and $\sqrt{f\left(\frac{h}{n}\right)}\to 1$ as $n\to\infty$, and we have the convergence of $\mathbb{E}\left[(\bar{J}_{n})^{2}\right]$ and $\mathbb{E}\left[\bar{J}_{n}J_{n}\right]$ by dominated convergence theorem. Thus, $\mathbb{E}\left[(\bar{J}_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right]\to 0$ as $n\to\infty$. By an identical argument, $\mathbb{E}\left[(J_{n}^{\beta})^{2}-\bar{J}_{n}^{\beta}J_{n}^{\beta}\right]\to 0$ as $n\to\infty$. This completes the proof of the convergence $\bar{J}_{n}-J_{n}$ to 0 in $L^{2}$, and hence the convergence $\bar{W}_{n}^{\beta}-W_{n}^{\beta}$ to 0 in $L^{1}$. ∎
3.0.1 Remark:
It follows from the proof that such convergence holds jointly for finitely many functions $f$ that satisfy the properties specified in the present paper. However, this proof strategy does not automatically extend to the uniform convergence in such space of functions. As a counterexample, take the sequence of functions $(f_{n})_{n\geq 1}$ defined by $f_{n}(x)=(1-x)^{n}$ for all $n\geq 1$.
As mentioned in the proof, we establish the $L^{2}$-convergence of $J_{n}$ as introduced in the proof of Theorem 3.3. Together with the observation that $n_{0}\geq 0$ may be chosen arbitrarily large (thereby making $K_{n}$ arbitrarily small in $L^{1}$), this implies that $(W_{n}^{\beta})_{n\geq 0}$ converges in $L^{1}$ to $W_{\infty}^{\beta}$ within the subcritical regime $\beta<\sqrt{2\log d}$. By identical lines of reasoning, the same convergence result holds for time-inhomogeneous environments, which is $\bar{J}_{n}$ appearing in the proof of Theorem 3.3.
Lemma 3.4.
Under the notation of Theorem 3.3’s proof, $J_{n}$ converges in $L^{2}$ for each fixed $n_{0}\geq 0$.
Proof.
We prove convergence by showing for each fixed $n_{0}\geq 0$, $J_{n}$ is a Cauchy sequence in $L^{2}$. For this, note that
$\mathbb{E}\left[(J_{n}-J_{m})^{2}\right]=\mathbb{E}\left[J_{n}^{2}\right]+\mathbb{E}\left[J_{m}^{2}\right]-2\mathbb{E}\left[J_{n}J_{m}\right].$
Since $J_{n}$ is uniformly bounded in $L^{2}$, it suffices to establish two inequalities: an upper bound for $\mathbb{E}\left[J_{n}^{2}\right]$ and a lower bound for $\mathbb{E}\left[J_{n}J_{m}\right]$:
$\limsup_{n\to\infty}\mathbb{E}\left[J_{n}^{2}\right]\leq\sum_{h=0}^{\infty}d^{-h}e^{\beta^{2}h}g^{\alpha}(h)$ (7)
and
$\liminf_{n,m\to\infty}\mathbb{E}\left[J_{n}J_{m}\right]\geq\sum_{h=0}^{\infty}d^{-h}e^{\beta^{2}h}g^{\alpha}(h),$ (8)
where for each $h\geq 0$ (corresponding to $C(\mathbf{x},\mathbf{y})=h$), $g^{\alpha}(h)$ is defined by:
$g^{\alpha}(h)=\tilde{\mathbb{P}}^{h}\left\{G^{\alpha}(\mathbf{x})\cap\tilde{G}^{\alpha}(\mathbf{x})\right\}:=\lim_{n\to\infty}\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap\tilde{G}_{n}^{\alpha}(\mathbf{x})\right\}.$
Lemma 3.5 guarantees this convergence is uniform for $C(\mathbf{x},\mathbf{y})=h<n_{1}$ and any fixed $n_{1}\geq 0$.
Our approach involves decomposing the series for $\mathbb{E}\left[J_{n}^{2}\right]$ and $\mathbb{E}\left[J_{n}J_{m}\right]$ into two parts: sums over pairs $(\mathbf{x},\mathbf{y})$ with $C(\mathbf{x},\mathbf{y})=h<n_{1}$, and their complements. As shown previously, the former dominates the contribution to $\mathbb{E}\left[J_{n}^{2}\right]$ for large $n_{1}$ (with the complement being negligible). Additionally, the probability of being in the good environment converges to $g^{\alpha}(h)$ uniformly for $h<n_{1}$.
For (7), we use the transformed measure $\tilde{\mathbb{P}}^{h}$ from the $L^{2}$-boundedness proof of Theorem 3.3. Fix $n_{1}\in\{n_{0},\ldots,n\}$ (eventually tending to infinity) and decompose the series by the number of differing steps between $\mathbf{x}$ and $\mathbf{y}$:
$\mathbb{E}\left[J_{n}^{2}\right]=\sum_{h=1}^{n_{1}}d^{-h}e^{\beta^{2}h}\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap\tilde{G}_{n}^{\alpha}(\mathbf{x})\right\}+\sum_{h=n_{1}+1}^{n}d^{-h}e^{\beta^{2}h}\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap\tilde{G}_{n}^{\alpha}(\mathbf{x})\right\}$
$\leq\sum_{h=1}^{n_{1}}d^{-h}e^{\beta^{2}h}\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap\tilde{G}_{n}^{\alpha}(\mathbf{x})\right\}+F^{\alpha}(n_{1}),$
where $F^{\alpha}(n_{1})$ can be made arbitrarily small (uniformly in $n$) by choosing $n_{1}$ sufficiently large. As in Theorem 3.3’s proof, $F^{\alpha}(n_{1})$ is bounded by the tail of the convergent series (5). Letting $n\to\infty$ and applying the dominated convergence theorem (with domination by the $L^{2}$-bound from (5)) yields (7).
For (8), consider $\mathbf{x}$ with $|\mathbf{x}|=m$, $\mathbf{y}$ with $|\mathbf{y}|=n$, and $C(\mathbf{x},\mathbf{y})=h\leq n_{1}\leq\min\{m,n\}$. Define the transformed measure for differing-length walks (where $\tilde{H}_{m}(\mathbf{x})$ denotes the Hamiltonian of length-$m$ branch $\mathbf{y}$):
$\frac{d\tilde{\mathbb{P}}^{h}}{d\mathbb{P}}=\frac{e^{\beta\left(H_{n}(\mathbf{x})+\tilde{H}_{m}(\mathbf{x})\right)}}{\mathbb{E}\left[e^{\beta\left(H_{n}(\mathbf{x})+\tilde{H}_{m}(\mathbf{x})\right)}\right]}.$
The argument mirrors that of (7): uniform convergence from Lemma 3.5 provides a lower bound when restricting to $C(\mathbf{x},\mathbf{y})<n_{1}$, with the complement becoming negligible as $n_{1}$ becomes large.
Combining (7) and (8) shows $J_{n}$ is a Cauchy sequence in $L^{2}$, hence convergent in $L^{2}$ for each fixed $n_{0}\geq 0$. ∎
The following lemma, which was used in the proof of the previous Lemma, establishes the uniform convergence of the events under the tilted probability:
Lemma 3.5.
Let $n_{1}>0$ be a fixed integer. For all $h\in\{0,\ldots,n_{1}\}$, the convergence
$\lim_{m,n\to\infty}\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap G_{m}^{\alpha}(\mathbf{y})\right\}=g^{\alpha}(h):=\mathbb{P}\left\{\tilde{G}^{\alpha,h}(\mathbf{x})\cap\tilde{G}^{\alpha,h}(\mathbf{y})\right\}$
uniformly for all $\mathbf{x}$ with $|\mathbf{x}|=n$, $\mathbf{y}$ with $|\mathbf{y}|=m$, and $C(\mathbf{x},\mathbf{y})=h$. Here, $\tilde{G}^{\alpha,h}$ is defined by
$\tilde{G}^{\alpha,h}(\mathbf{x})=\left\{\forall T\geq n_{0},\,\tilde{H}^{h}_{T}(\mathbf{x})<\alpha T\right\},$
where $\tilde{H}^{h}_{T}$ denotes the tilted Hamiltonian (i.e., the original Hamiltonian $H_{T}$ under the transformed measure $\tilde{\mathbb{P}}^{h}$). Moreover, the same convergence (to the same limit) holds for time-inhomogeneous weights:
$\lim_{m,n\to\infty}\tilde{\mathbb{P}}^{h}\left\{\bar{G}_{n}^{\alpha}(\mathbf{x})\cap\bar{G}_{m}^{\alpha}(\mathbf{y})\right\}=g^{\alpha}(h):=\mathbb{P}\left\{\tilde{G}^{\alpha,h}(\mathbf{x})\cap\tilde{G}^{\alpha,h}(\mathbf{y})\right\}$
Proof.
For each $\mathbf{x}$ with $|\mathbf{x}|=n$ and $T\in\{0,\ldots,n\}$, let $E^{\alpha}(T)$ to be the event $E^{\alpha}_{T}(\mathbf{x})=\left\{H_{T}(\mathbf{x})<\alpha T\right\}$ and define $\tilde{E}^{\alpha,h}_{T}(\mathbf{x})$ analogously for $\tilde{H}^{h}_{T}(\mathbf{x})$. By definition, $\tilde{G}^{\alpha,h}=\bigcap_{T=n_{0}}^{\infty}\tilde{E}^{\alpha,h}_{T}(\mathbf{x})$.
We first establish the upper bound:
$\limsup_{m,n\to\infty}\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap G_{m}^{\alpha}(\mathbf{y})\right\}\leq g^{\alpha}(h).$ (9)
To see this, observe that for any fixed $n_{2}>n_{0}$, when $m,n>n_{2}$, we have
$\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap G_{m}^{\alpha}(\mathbf{y})\right\}\leq\mathbb{P}\left\{\bigcap_{T=n_{0}}^{n_{2}}\tilde{E}^{h}_{T}(\mathbf{x})\cap\tilde{E}^{h}_{T}(\mathbf{y})\right\}.$
Since $n_{2}>n_{0}$ is arbitrary, taking $n_{2}\to\infty$ gives
$\limsup_{m,n\to\infty}\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap G_{m}^{\alpha}(\mathbf{y})\right\}\leq\tilde{\mathbb{P}}^{h}\left\{\bigcap_{T=n_{0}}^{\infty}E^{\alpha,h}_{T}(\mathbf{x})\cap E^{\alpha,h}_{T}(\mathbf{y})\right\}=\tilde{\mathbb{P}}^{h}\left\{G^{\alpha}(\mathbf{x})\cap G^{\alpha}(\mathbf{y})\right\}=g^{\alpha}(h),$
confirming (9).
For the lower bound, we use a similar argument to that in Theorem 3.3. For any fixed $\epsilon>0$, there exists $n_{2}$ (independent of $m,n$) such that for all $m,n>0$,
$\tilde{\mathbb{P}}^{h}\left\{\bigcup_{T=n_{2}}^{\max\{m,n\}}\left(H_{T}(\mathbf{x})\geq\alpha T\right)\right\}<\epsilon,$
with the empty event assigned probability 0. The key is to observe that the constraint $H_{n}(\mathbf{x})<\alpha n$ defining the event $E_{n}(\mathbf{x})$ is essentially guaranteed for large $n$ under $\tilde{\mathbb{P}}^{h}$: when $C(\mathbf{x},\mathbf{y})<n_{1}$ so the ranches $\mathbf{x}$ and $\mathbf{y}$ are well separated (note that this would be false however if $\mathbf{x}$ and $\mathbf{y}$ would not be separated: for instance if $\mathbf{x}=\mathbf{y}$, then $\tilde{H}_{n}(\mathbf{x})$ would be of the order of $2\beta n$, which is larger than $\alpha n$.) As a result, this follows directly from the proof of Theorem 3.3 (where $\tilde{\mathbb{P}}\{G_{n}^{\alpha}(\mathbf{x})\}\geq 1-\epsilon(n_{0})$) by replacing $n_{0}$ with $n_{2}$.
Consequently, the lower bound holds:
$\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap G_{m}^{\alpha}(\mathbf{y})\right\}\geq\mathbb{P}\left\{\bigcap_{T=n_{0}}^{n_{2}}\tilde{E}^{\alpha,h}_{T}(\mathbf{x})\cap\tilde{E}^{\alpha,h}_{T}(\mathbf{y})\right\}-2\epsilon.$ (10)
Letting $m,n\to\infty$ yields
$\liminf_{m,n\to\infty}\tilde{\mathbb{P}}^{h}\left\{G_{n}^{\alpha}(\mathbf{x})\cap G_{m}^{\alpha}(\mathbf{y})\right\}\geq\mathbb{P}\left\{\bigcap_{T=n_{0}}^{n_{2}}\tilde{E}^{\alpha,h}_{T}(\mathbf{x})\cap\tilde{E}^{\alpha,h}_{T}(\mathbf{y})\right\}-2\epsilon$
$\geq\mathbb{P}\left\{\bigcap_{T=n_{0}}^{\infty}\tilde{E}^{\alpha,h}_{T}(\mathbf{x})\cap\tilde{E}^{\alpha,h}_{T}(\mathbf{y})\right\}-2\epsilon$
Since $\epsilon>0$ is arbitrary, this establishes (10). Combining (9) and (10) completes the proof, with uniformity in $C(\mathbf{x},\mathbf{y})\leq n_{1}$ preserved throughout.
The uniform convergence to the same limit also holds for time-inhomogeneous weights. When $f$ is decreasing, for fixed $n_{2}\geq 0$, when $n$ is large enough (such that $\frac{n_{2}}{n}\to 0$ and $f\left(\frac{h}{n}\right)\to 1$ for all $h\in\{0,\dots,n_{2}\}$), the probability of the following finite intersection of events converges:
$\mathbb{P}\left\{\bigcap_{T=n_{0}}^{n_{2}}\tilde{\bar{E}}^{h}_{T}(\mathbf{x})\cap\tilde{\bar{E}}^{h}_{T}(\mathbf{y})\right\}\to\mathbb{P}\left\{\bigcap_{T=n_{0}}^{n_{2}}\tilde{E}^{h}_{T}(\mathbf{x})\cap\tilde{E}^{h}_{T}(\mathbf{y})\right\},$
as $n\to\infty$, by continuity of Gaussian cumulative density function. Here, $\tilde{\bar{E}}^{h}_{T}(\mathbf{x})$ is defined as $\tilde{\bar{E}}^{h}_{T}(\mathbf{x}):=\left\{\bar{H}_{T}<\alpha\sum_{i=1}^{T}f\left(\frac{i}{n}\right)\right\}$.
As a result, we obtain the same upper bound since $n_{2}\geq 0$ can be chosen arbitrarily. The lower bound, meanwhile, can be derived via the same argument to that used for the time-homogeneous case in the proof above.∎
The following proposition establishes the upper bound for $\bar{\beta}_{c}$, i.e. when $\beta>\beta_{c}$, $\bar{W}_{n}^{\beta}$ converges to 0 in probability. Together with Theorem 3.3, this confirms that the normalized partition function $\bar{W}_{n}^{\beta}$ exhibits distinct phase transitions in its limit in probability. Furthermore, the critical parameter $\bar{\beta}_{c}$ is defined by:
$\bar{\beta}_{c}:=\sup\left\{\beta\geq 0\mid\mathbb{P}\{\bar{W}_{\infty}^{\beta}=0\}>0\right\}$
exists and coincides with that for homogeneous weights, i.e., $\beta_{c}=\bar{\beta}_{c}$. This result underscores the insensitivity to environmental variations, indicating that the macroscopic behavior of the partition function is dominated by the ”early generations” of the branching structure, rather than being affected by prescribed variations in the branch weights.
Theorem 3.6.
Let $\beta>\sqrt{2\log d}$ and let $f:[0,1]\to[0,1]$ be a continuous decreasing function satisfying $f(0)=1$ and $f(1)=0$. Then the normalized partition function $\bar{W}_{n}^{\beta}$ converges in probability to the random variable $\bar{W}_{\infty}^{\beta}\equiv 0$ as $n\to\infty$. As a consequence, $\bar{\beta}_{c}:=\sup\left\{\beta\geq 0\mid\mathbb{P}\{\bar{W}_{\infty}^{\beta}=0\}>0\right\}$ exists and coincides with that for homogeneous weights, i.e., $\beta_{c}=\bar{\beta}_{c}$.
Proof.
We fix a constant $\alpha\in(0,1)$ and it suffices to show that $\mathbb{E}\left[(\bar{W}_{n}^{\beta})^{\alpha}\right]\to 0$, as this implies $\bar{W}_{n}^{\beta}$ converges in probability to $\bar{W}_{\infty}^{\beta}\equiv 0$.
First, note that by the independence of weight increments, the partition function $\bar{W}_{n}^{\beta}$ can be decomposed as:
$\bar{W}_{n}^{\beta}=\sum_{|\mathbf{x}|=n_{1}}d^{-n_{1}}e^{-\frac{1}{2}\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+\beta\bar{H}_{n_{1}}(\mathbf{x})}\sum_{\mathbf{y}\in D^{\mathbf{x}}_{n-n_{1}}}d^{-(n-n_{1})}e^{-\frac{1}{2}\beta^{2}\sum_{i=n_{1}+1}^{n}f\left(\frac{i}{n}\right)+\beta\bar{H}_{n_{1}\to n}(\mathbf{y})},$
where $D^{\mathbf{x}}_{n-n_{1}}$ denotes the set of offspring in the sub-GW tree rooted at a branch $\mathbf{|x|}=n_{1}$ with $n-n_{1}$ generations, and $\bar{H}_{n_{1}\to n}(\mathbf{x})=\sum_{i=n_{1}}^{n}\sqrt{f\left(\frac{i}{n}\right)}w(x_{i})$ is the Hamiltonian from generation $n_{1}$ to $n$.
We define a transformed (random) probability measure $\bar{\mathbb{P}}^{n_{1}}$ by (where $\mathbf{P}$ is the law of the BRWs):
$\frac{\mathrm{d}\bar{\mathbb{P}}^{n_{1}}}{\mathrm{d}\mathbf{P}}=\frac{e^{-\frac{1}{2}\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+\beta\bar{H}_{n_{1}}(\mathbf{x})}}{\bar{W}_{n}^{n_{1}}}$ (11)
Let $\mathbf{E}_{n_{1}}$ denote the expectation under $\bar{\mathbb{P}}^{n_{1}}$, and abbreviate the decomposed partition function by:
$\bar{W}_{n}^{n_{1}}=\sum_{|\mathbf{x}|=n_{1}}d^{-n_{1}}e^{-\frac{1}{2}\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+\beta\bar{H}_{n_{1}}(\mathbf{x})}$
and
$\tilde{\bar{W}}_{n}^{n_{1}}=\sum_{\mathbf{y}\in D^{\mathbf{x}}_{n-n_{1}}}d^{-(n-n_{1})}e^{-\frac{1}{2}\beta^{2}\sum_{i=n_{1}+1}^{n}f\left(\frac{i}{n}\right)+\beta\bar{H}_{n_{1}\to n}(\mathbf{y})}.$
For any function $F:D_{n_{1}}\to\mathbb{R}$ (recall that $D_{n_{1}}$ is the set of branches $|\mathbf{x}|=n$), we have
$\mathbf{E}_{n_{1}}[F(\mathbf{x})]=\frac{1}{\bar{W}_{n}^{n_{1}}}\sum_{|\mathbf{x}|=n_{1}}\frac{1}{d^{n_{1}}}e^{\beta\bar{H}_{n_{1}}(\mathbf{x})-\frac{1}{2}\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)}F(\mathbf{x}),$
which is $\mathcal{F}_{n_{1}}$-measurable (recall that $\mathcal{F}_{n_{1}}$ is the $\sigma$-field generated by the GW tree and the weights up to generation $n_{1}$).
It follows that
$\mathbb{E}\left[(\bar{W}_{n}^{\beta})^{\alpha}\right]=\mathbb{E}\left[(\bar{W}_{n}^{n_{1}})^{\alpha}\left(\mathbf{E}_{n_{1}}\left[\tilde{\bar{W}}_{n}^{n_{1}}\right]\right)^{\alpha}\right].$
By the Cauchy-Schwarz inequality, this implies
$\mathbb{E}\left[(\bar{W}_{n}^{\beta})^{\alpha}\right]\leq\mathbb{E}\left[(\bar{W}_{n}^{n_{1}})^{2\alpha}\right]^{\frac{1}{2}}\mathbb{E}\left[\left(\mathbf{E}_{n_{1}}\left[\tilde{\bar{W}}_{n}^{n_{1}}\right]\right)^{2\alpha}\right]^{\frac{1}{2}}.$ (12)
To bound the second term on the right hand side in (12), for $0<\alpha<\frac{1}{2}$, Jensen’s inequality implies that:
$\mathbb{E}\left[\left(\mathbf{E}_{n_{1}}\left[\tilde{\bar{W}}_{n}^{n_{1}}\right]\right)^{2\alpha}\right]\leq\mathbb{E}\left[\mathbf{E}_{n_{1}}\left[\tilde{\bar{W}}_{n}^{n_{1}}\right]\right]^{2\alpha}$
By the independence of $\tilde{\bar{W}}_{n}^{n_{1}}$ and the $\sigma$-field $\mathcal{F}_{t}$ (where the tilted probability measure $\bar{\mathbb{P}}^{n_{1}}$ is $\mathcal{F}_{t}$-measurable), we have the expectation are the same under two probability measures:
$\mathbb{E}\left[\mathbf{E}_{n_{1}}\left[\tilde{\bar{W}}_{n}^{n_{1}}\right]\right]^{2\alpha}=\mathbb{E}\left[\mathbb{E}\left[\tilde{\bar{W}}_{n}^{n_{1}}\right]\right]^{2\alpha}=1$
To bound $\mathbb{E}\left[(\bar{W}_{n}^{n_{1}})^{2\alpha}\right]$, we will show it by computing its fractional moment:
$\mathbb{E}\left[(\bar{W}_{n}^{n_{1}})^{2\alpha}\right]\leq d^{n_{1}}\mathbb{E}\left[d^{-2\alpha n}e^{-\alpha\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+2\alpha\beta\bar{H}_{n_{1}}(\mathbf{x})}\right],$
which simplifies to:
$\leq d^{(1-2\alpha)n_{1}}e^{-\alpha\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)}\mathbb{E}\left[e^{2\alpha\beta\bar{H}_{n_{1}}(\mathbf{x})}\right]$
and further to:
$\leq d^{(1-2\alpha)n_{1}}e^{-\alpha\beta^{2}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)}e^{2\alpha^{2}\beta^{2}n_{1}}=e^{(\log d(1-2\alpha)-\frac{\alpha\beta^{2}}{n_{1}}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+2\alpha^{2}\beta^{2})n_{1}}.$ (13)
Finally, for any fixed $n_{1}>0$ and $\beta>\sqrt{2\log d}$, we have when letting $n\to\infty$:
$(1-2\alpha)\log d-\frac{\alpha\beta^{2}}{n_{1}}\sum_{i=1}^{n_{1}}f\left(\frac{i}{n}\right)+2\alpha^{2}\beta^{2}\to(1-2\alpha)\log d-\alpha\beta^{2}+2\alpha^{2}\beta^{2}$
This convergence holds because $f\left(\frac{i}{n}\right)$ can be arbitrarily close to $1$ for all $i<n_{1}$ when $n$ becomes sufficiently large. Subsequently, by letting $\alpha$ approach $\frac{1}{2}$ arbitrarily close, we have $\log d\cdot(1-2\alpha)-\alpha\beta^{2}+2\alpha^{2}\beta^{2}<0$. Moreover, the right-hand side of (13) can be made arbitrarily small by choosing $n_{1}$ sufficiently large. Combining with the inequality (12), we conclude that $\bar{W}_{n}^{\beta}\to 0$ in probability when $\beta>\beta_{c}=\sqrt{2\log d}$. ∎
Using the strategy of the proof of these results, we end up this section by providing a novel proof for Theorem 2.1. Since the proof follows from the similar lines with that of Theorem 3.3, we prefer not to state all the details. Indeed, we can see that the proof is much simpler compared to the proof of Theorem 3.3: since $(W_{n}^{\beta})_{n\geq 0}$ defines a martingale, to show $L^{1}$ convergence it is suffice to show that it is uniformly integrable (i.e. $J_{n}$ is uniformly bounded in $L^{2}$ for each fixed $n_{0}>0$). The upper bound follows from the same argument as the last part of the proof of Proposition 3.6: for $\alpha\in(0,1)$ we have:
$\mathbb{E}\left[(W_{n}^{\beta})^{\alpha}\right]\leq d^{n}\mathbb{E}\left[d^{-\alpha n}e^{-\frac{1}{2}\alpha\beta^{2}n+\alpha\beta H_{n}(\mathbf{x})}\right]=e^{n\left((1-\alpha)\log d-\frac{1}{2}\alpha\beta^{2}+\frac{1}{2}\alpha^{2}\beta^{2}\right)}.$
We can see that if $\beta$ satisfies $\log d-\frac{1}{2}\alpha\beta^{2}<0$, then $\mathbb{E}\left[(W_{n}^{\beta})^{\alpha}\right]\to 0$ as $n\to\infty$, and we may choose $\alpha$ arbitrarily close to $1$ so that it holds for any $\beta>\sqrt{2\log d}$.
3.1 Behavior of Partition Function at Criticality $\bar{\beta}_{c}$
For a long time—over two decades after the establishment of the phase transition of Biggins martingale—the exact behavior of the time-homogeneous partition function $W_{n}^{\beta}$ at the critical value $\beta=\beta_{c}$ remained unknown, except that it lies in the strong disorder regime (i.e. $W_{n}^{\beta}$ converges to 0 almost surely). Naturally, this raises the question of the decay rate of $\bar{W}_{n}^{\beta}$ at criticality. Since $\beta_{c}=\bar{\beta_{c}}$ (as shown in Theorem 3.3), we will henceforth denote both by $\beta_{c}$ for the remainder of this section.
The problem of determining the exact asymptotic behavior at criticality was resolved in the influential work of Hu and Shi [12], who established the following key result (Theorem 1.3 in [12]):
$W_{n}^{\beta_{c}}=n^{-\frac{1}{2}+o(1)}\quad\text{a.s.}$
Notably, this implies that even though $\beta=\beta_{c}$ lies in the strong disorder regime, the partition function decays only polynomially. The next theorem suggests that this asymptotic behavior extends to time-inhomogeneous weights as well, and that the universality principle extends to the critical regime (though convergence is now almost sure).
Theorem 3.7.
(Asymptotics of partition function at criticality) At the critical parameter $\beta_{c}$, we have $\bar{W}_{n}^{\beta_{c}}=n^{-\frac{1}{2}+o(1)}$ almost surely.
Proof.
Our main goal in this step is see that it is suffice to show that the same asymptotics holds for all fractional moments $\alpha\in(0,1)$. The arguments here are fairly standard, and variations of these arguments are also employed in related works [12] and [2]:
Lemma 3.8.
(Upper bounds) Let $(a_{n})_{n\in\mathbb{N}}$ be a sequence of real numbers such that $|a_{n}|\to\infty$ as $n\to\infty$. Suppose that for every $\alpha\in(0,1)$, we have:
$\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]=e^{\alpha a_{n}(1+o(1))}.$
Then $\bar{W}_{n}^{\beta_{c}}\leq e^{a_{n}(1+o(1))}$ in probability as $n\to\infty$. For our purposes, we set $a_{n}=-\frac{1}{2}\log n$. Moreover, if for any $\epsilon>0$, $\sum_{n\geq 1}e^{-\epsilon|a_{n}|}<\infty$, then $\bar{W}_{n}^{\beta_{c}}\leq e^{a_{n}(1+o(1))}$ almost surely.
Proof.
Fix $\epsilon>0$ and let $\alpha\in(0,1)$. By the given assumption, $\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]\leq e^{\alpha a_{n}+\epsilon\alpha|a_{n}|/2}$ for all sufficiently large $n$. Applying Markov’s inequality, we obtain:
$\mathbb{P}(\bar{W}_{n}^{\beta_{c}}>e^{a_{n}+\epsilon|a_{n}|})=\mathbb{P}((\bar{W}_{n}^{\beta_{c}})^{\alpha}>e^{\alpha a_{n}+\epsilon\alpha|a_{n}|})$ $\leq e^{-\alpha a_{n}-\epsilon\alpha|a_{n}|}\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]$
$\leq e^{-\epsilon\alpha|a_{n}|/2}$
By the assumption that $|a_{n}|\to\infty$, it follows that $\bar{W}_{n}^{\beta_{c}}\leq e^{a_{n}(1+o(1))}$ in probability. The almost sure convergence of the statement follows from Borel-Cantelli lemma. ∎
Lemma 3.9 (Lower bounds).
Let $(a_{n})_{n\in\mathbb{N}}$ be a sequence satisfying $(\log n)^{\frac{1}{2}}\ll|a_{n}|\ll n$. Assume that for all $\alpha\in(0,1)$, we have:
$\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]=e^{\alpha a_{n}(1+o(1))}.$
Then almost surely,
$\bar{W}_{n}^{\beta_{c}}\geq e^{a_{n}(1+o(1))}.$
Proof.
We abbreviate the notation by defining the ”partial free energy” for the time-inhomogeneous weight as:
$\bar{V}_{\tau_{n}}^{n}(w)=\beta_{c}\bar{H}_{n}^{\tau_{n}}(w)-\frac{1}{2}\beta_{c}^{2}\sum_{i=1}^{\tau_{n}}f(\frac{i}{n})-\tau_{n}\log d$
Fix $\epsilon>0$, and we will show that the probability that $\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]$ is beyond the rate $e^{\alpha a_{n}\pm\epsilon\alpha|a_{n}|}$ tends to 0 as $n\to\infty$. By the given assumption of the asymptotic rate of fractional moment, for any $\alpha\in(0,\frac{1}{2})$, we have:
$\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]\geq e^{\alpha a_{n}-\frac{\epsilon}{4}\alpha|a_{n}|}$
and
$\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{2\alpha}]\leq e^{2\alpha a_{n}+\frac{\epsilon}{4}|a_{n}|}$
for all sufficiently large $n$. Applying the Paley-Zygmund inequality, we derive:
$\mathbb{P}\left(\bar{W}_{n}^{\beta_{c}}>e^{a_{n}-\epsilon|a_{n}|}\right)$ $\geq\Big(1-\frac{e^{\alpha(a_{n}-\epsilon|a_{n}|)}}{\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]}\Big)^{2}\frac{\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{\alpha}]^{2}}{\mathbb{E}[(\bar{W}_{n}^{\beta_{c}})^{2\alpha}]}$
$\geq\big(1-e^{-3\epsilon\alpha|a_{n}|/4}\big)^{2}e^{-\frac{3\epsilon}{4}|a_{n}|}\geq e^{-\epsilon|a_{n}|}\,$, (14)
for all sufficiently large $n$. Now define $\tau_{n}$ to be:
$\tau_{n}=\lceil\frac{2\epsilon|a_{n}|}{\log d}\rceil$
so that $\tau_{n}<n$ for all sufficiently large $n$ and $\tau_{n}\ll n$. We then have the decomposition:
$\bar{W}_{n}^{\beta_{c}}$ $=\sum_{|w|=\tau_{n}}\frac{1}{d^{\tau_{n}}}e^{\beta_{c}\bar{H}_{\tau_{n}}(w)-\frac{1}{2}\beta_{c}^{2}\sum_{i=1}^{\tau_{n}}f(\frac{i}{n})}\!\!\!\!\sum_{\begin{subarray}{c}v\in D_{n-\tau_{n}}(w)\end{subarray}}\frac{1}{d^{n-\tau_{n}}}e^{\beta_{c}\bar{H}_{\tau_{n}\to n}(w)-\frac{1}{2}\beta_{c}^{2}\sum_{i=\tau_{n}+1}^{n}f(\frac{i}{n})}$
$\geq\exp\big\{\max_{|w|=\tau_{n}}\bar{V}_{\tau_{n}}^{n}(w)\big\}\sum_{|w|=\tau_{n}}\!\!\!\ \sum_{\begin{subarray}{c}v\in D_{n-\tau_{n}}(w)\end{subarray}}e^{\beta_{c}\bar{H}_{\tau_{n}\to n}(w)-\frac{1}{2}\beta_{c}^{2}\sum_{i=\tau_{n}+1}^{n}f(\frac{i}{n})}$
We denote the rightmost sum by $Y_{n-\tau_{n}}(w)$. The above inequality then yields the following estimate:
$\mathbb{P}\Big(\bar{W}_{n}^{\beta_{c}}\leq e^{a_{n}-\epsilon|a_{n}|}\exp\{\max_{|v|=\tau_{n}}\bar{V}_{\tau_{n}}^{n}(v)\}\Big)$ $\leq\mathbb{P}\Big(\sum_{|w|=\tau_{n}}Y_{n-\tau_{n}}(v)\leq e^{a_{n}-\epsilon|a_{n}|}\Big)$
$\leq\mathbb{P}\left(Y_{n-\tau_{n}}(v)\leq e^{a_{n}-\epsilon|a_{n}|}\right)^{d^{\tau_{n}}}$
Equation (14) implies that this expression is bounded above by $\exp\{-e^{\epsilon|a_{n}|}\}$ (since the fluctuations of the increments are decreasing). Therefore, by the assumption that $|a_{n}|\gg(\log n)^{\frac{1}{2}}$, these probabilities are summable. By the Borel-Cantelli lemma, it follows that almost surely
$Y_{n}\geq e^{a_{n}-\epsilon|a_{n}|}\exp\big\{\max_{|v|=\tau_{n}}\bar{V}_{\tau_{n}}^{n}(v)\big\}\,$, (15)
for all sufficiently large $n$.
However, it is well known that there exists an explicit constant $C>0$ such that:
$\frac{1}{\tau_{n}}\max_{|v|=\tau_{n}}\{\bar{V}_{\tau_{n}}^{n}(v)\}\to C$
almost surely (this maximum corresponds to the position of the rightmost particle in the time-homogeneous branching random walk system. In this case, since $\tau_{n}\ll n$, this result also holds for the first $\tau_{n}$ steps of the time-inhomogeneous environment). Thus,
$\exp\big\{\max_{|v|=\tau_{n}}\{\bar{V}_{\tau_{n}}^{n}(v)\}\big\}=e^{C^{*}\epsilon|a_{n}|(1+o(1))}$
for some constant $C^{*}>0$ (independent of $\epsilon$). Hence, almost surely, $Y_{n}\geq e^{a_{n}(1+o(1))}$. ∎
To conclude the proof, we apply the following classical result in GMC theory, which is the so-called Kahane’s concentration inequality (this inequality originated in [13]; see [16] for a proof):
Lemma 3.10.
Let $F$ be a concave function such that:
$\forall x\in\mathbb{R}_{+},|F(x)|\leq M\left(1+|x|^{a}\right)$
for some positive constants $M,a$, and let $\sigma$ be a Radon measure on the Borel subsets of $\mathbb{R}^{d}$. Given a bounded Borel set $A$, let $(X_{r})_{r\in A},(Y_{r})_{r\in A}$ be two continuous centered Gaussian processes with continuous covariance kernels $K_{X}$ and $K_{Y}$ such that:
$\forall u,v\in A,K_{X}(u,v)\leq K_{Y}(u,v)$
Then we have:
$\mathbb{E}\left[F\left(\int_{A}e^{X_{r}-\frac{1}{2}\mathbb{E}[X_{r}^{2}]}\sigma(dr)\right)\right]\geq\mathbb{E}\left[F\left(\int_{A}e^{Y_{r}-\frac{1}{2}\mathbb{E}[Y_{r}^{2}]}\sigma(dr)\right)\right]$
In this case, we may couple the law of the G-W tree for both time-homogeneous and time-inhomogeneous BRWs such that they have the same offspring distribution, and embed each realization of the G-W tree into the unit interval (i.e., $A=[0,1]$; see Section 2 for the detailed construction of such an embedding), and let the Borel measure $\sigma$ be the point mass associated to each realization. Note that for any $\alpha<1$, $F(x)=x^{\alpha}$ satisfies the conditions in the lemma.
The upper bound of $\bar{W}_{n}^{\beta_{c}}$ follows by a similar argument as in Proposition 3.6, except that Hölder’s inequality is used here instead of the Cauchy-Schwarz inequality. We need to show that for any $\epsilon^{\prime}>0$, we have $\mathbb{E}[(\bar{W}_{n}^{\beta})^{\alpha}]\leq n^{-\frac{1}{2}+\epsilon^{\prime}}$. For all $p,q>1$ satisfying $1/p+1/q=1$, and recalling the expectation $\mathbf{E}_{n_{1}}$ under the (random) measure $\bar{\mathbb{P}}^{n_{1}}$ defined in Section 11, we have:
$\mathbb{E}\left[(\bar{W}_{n}^{\beta})^{\alpha}\right]\leq\mathbb{E}\left[(\bar{W}_{n}^{n^{1-\epsilon}})^{p\alpha}\right]^{1/p}\mathbb{E}\left[\left(\mathbf{E}_{n^{1-\epsilon}}\left[\tilde{\bar{W}}_{n}^{n^{1-\epsilon}}\right]\right)^{q\alpha}\right]^{1/q}$
Note that for any $q>1$, the second term on the right-hand side of the above inequality equals 1. As a result, for any fixed $\alpha>0$, $n_{1}>0$ and $\epsilon>0$, we may let $p$ be arbitrarily close to 1 so that $p\alpha<1$. Since $f$ is continuous at 0, for any fixed $\epsilon>0$, when $n$ is large enough, there exists $\delta>0$ such that $\sum_{i=1}^{n^{1-\epsilon}}\left(f(\frac{i}{n})-1\right)<(1-\delta)n$. By Kahane’s concentration inequality (Lemma 3.10), we can see that:
$\mathbb{E}[(W_{(1-\delta)n}^{\beta})^{p\alpha}]\geq\mathbb{E}\left[(\bar{W}_{n}^{n^{1-\epsilon}})^{p\alpha}\right]$
As a result, since the asymptotic behavior of $W_{n}^{\beta_{c}}$ is $n^{-\frac{1}{2}+o(1)}$, we have (by the dominated convergence theorem) the convergence of the fractional moment:
$\mathbb{E}\left[(\bar{W}_{n}^{n^{1-\epsilon}})^{p\alpha}\right]\leq((1-\delta)n)^{-\frac{1}{2p}+o(1)}\leq n^{-\frac{1}{2p}(1-\delta)+o(1)}$
Then we can derive the upper bound for the fractional moment:
$\limsup_{n\to\infty}\mathbb{E}\left[(\bar{W}_{n}^{\beta})^{\alpha}\right]\leq n^{-\frac{1}{2p}(1-\delta)+o(1)}$
which holds for any $p>1$ and $\epsilon>0$. Letting $p\to 1$ and using the fact that $\delta$ can be made arbitrarily small by taking $\epsilon$ sufficiently small yields the desired upper bound.
Similarly, the lower bound is obtained via a direct comparison of the fractional moments of the two partition functions:
$\mathbb{E}\left[\left(\bar{W}_{n}^{\beta}\right)^{\alpha}\right]\geq\mathbb{E}\left[\left(W_{n}^{\beta}\right)^{\alpha}\right]$
This finishes the proof of the asymptotic behavior of $\bar{W}_{n}^{\beta}$ at $\beta=\bar{\beta}_{c}$. ∎
4 Further Directions
Here, we briefly contextualize our established results within the landscape of related current research areas (including BRWs, directed polymers and GMC) and outline potential directions for future work:
4.1 Behavior of Partition Function at Criticality $\bar{\beta}_{c}$
The derivative martingale is another key martingale associated with branching random walks and plays a central role in the study of their extreme values (see Section 5.4 of [17] for more details). At the critical parameter $\beta=\beta_{c}$, Biggins and Kyprianou [5] showed that the derivative martingale (as the name suggests, by taking ”derivative” with respect to $\beta$ of $W_{n}^{\beta}$ at $\beta_{c}$), defined by:
$D_{n}:=\sum_{|\mathbf{x}|=n}\left(\beta_{c}H_{n}(\mathbf{x})-\frac{\beta_{c}^{2}n}{2}\right)e^{\beta_{c}H_{n}(\mathbf{x})-\frac{\beta_{c}^{2}n}{2}}$
converges almost surely to a non-degenerate limit $D_{\infty}>0$. Furthermore, Elie and Shi [1] established that $D_{\infty}$ has the same law as a constant multiple of the limit in probability of $n^{\frac{1}{2}}W_{n}^{\beta_{c}}$ (the constant depends only on the law of the environment and the Galton-Watson tree):
Theorem 4.1 ([1], Theorem 1.1).
We have the Seneta–Heyde scaling for Derivative martingale $D_{\infty}$:
$\lim_{n\to\infty}n^{\frac{1}{2}}W_{n}^{\beta_{c}}=\left(\frac{2}{\pi\sigma^{2}}\right)^{\frac{1}{2}}D_{\infty}$
where $\sigma^{2}$ depends on the law of environment and G-W tree only.
While one may conjecture that a similar scaling result holds in time-inhomogeneous weights, it is worth mentioning the difference between the model in question and the time-homogeneous case. Either in the literature of critical GMC or in BRWs, the behavior of the partition function at criticality is governed by the extreme value (of the Gaussian field/BRWs). As we mentioned before, in the sub-critical regime ($\gamma<\sqrt{2d}$), the GMC measure is almost surely supported on $\gamma$-thick points. At the critical parameter $\gamma=\sqrt{2d}$, it corresponds to the maximal possible thickness:
$\lim_{\epsilon\to 0}\sup_{x\in D}X_{\epsilon}(x)/|\log\epsilon|\to\sqrt{2d}\text{ in probability}$
In the literature of BRWs, it was shown in [14] that when the variance is decreasing, the extreme value of BRWs with decreasing variance exhibits different asymptotic behavior from their constant variance counterparts. However, the asymptotics of the extreme value of the time-inhomogeneous BRWs studied in the present paper remains an open problem, which would be interesting to explore in the future.
4.2 Relationship with Two-Dimensional Directed Polymers
There is an analogy between two-dimensional directed polymers and branching random walks with decreasing variance, and the former serves as one of the primary motivation for the later model discussed in the current paper.
Directed polymers in random environment describe the behavior of a long directed chain of monomers in presence of random impurities, see [8] for a comprehensive introduction to this area. In this setting, the trajectory of the polymer is given by a nearest-neighbor path $(S_{n})_{n\in\mathbb{N}}$ on two-dimensional lattice, while the impurities (also called the environment) are given by i.i.d. centered random variables $\omega(n,x)$, $n\in\mathbb{N},x\in\mathbb{Z}^{d}$ of variance one satisfying the free energy $\lambda(\beta)=\log\mathbb{E}[e^{\beta\omega(n,x)}]\in(0,\infty)$ for all $\beta>0$. The polymer measure on the path space is then defined by:
$\text{d}P_{N,\beta}^{\omega}(S)=\frac{e^{\beta\sum_{n=1}^{N}\left(\omega(n,S_{n})-\lambda(\beta)\right)}}{Z_{N}(\beta)}\text{d}P(S),$ (16)
where $\beta>0$ is the inverse temperature, $P=P_{0}$ with $(P_{x},E_{x})$ the probability measure and expectation of the simple random walk on $\mathbb{Z}^{d}$ started at point $x\in\mathbb{Z}^{d}$, and $Z_{N}(\beta)=E[e^{\beta\sum_{n=1}^{N}\left(\omega(n,S_{n})-\lambda(\beta)\right)}]$ is the (normalized) partition function.
For $x\in\mathbb{Z}^{2}$, define $W_{N}(x,\beta)=\mathbb{E}_{x}\left[e^{\sum_{n=1}^{N}\left(\beta\omega(n,S_{n})-\lambda(\beta)\right)}\right]$ to be the partition function for paths start at $x$. In the siminal paper [6], Caravenna, Sun and Zygouras established the convergence of the spatial behavior of two-dimensional partition functions:
Theorem 4.2.
([6], Theorem 2.5) Assume $d=2$ and set $\beta_{N}$ defined above. For all $\hat{\beta}\geq 1$ one has $W_{N}(x,\beta_{N})$ converges to 0 in probability as $N\to\infty$, while the log-partition function satisfies a central limit theorem in the weak disorder regime ($\hat{\beta}<1$):
$\forall\hat{\beta}\in(0,1),\log W_{N}(x,\beta_{N})\to^{(d)}\mathcal{N}(-\frac{\lambda^{2}}{2},\lambda^{2}),\text{ where }\lambda^{2}:=\log(\frac{1}{1-\hat{\beta}^{2}})$
In [9], the authors showed that when $\omega(n,x)$ are i.i.d. standard Gaussian (i.e. having law $\mathcal{N}(0,1)$), where $\lambda(\beta)=\frac{1}{2}\beta^{2}$, the normalized partition function for two-dimensional directed polymers at $x\in\mathbb{Z}^{2}$:
$h_{N}(x):=\sqrt{\log N}\left(\log W_{N}(x\sqrt{N},\beta_{N})-\mathbb{E}[\log W_{N}(x\sqrt{N},\beta_{N})]\right)$
can be approximated by $M_{N}(x)=\sum_{k=1}^{M}X_{k}(x)$ for some $M>0$ (eventually large), where $X_{k}(x)$ is defined by the normalized ”$k$-th step” partition function:
$X_{k}(x):=\sqrt{\log N}\left(\log Z_{N}(x\sqrt{N},\beta_{N})-\mathbb{E}[\log Z_{N}(x\sqrt{N},\beta_{N})]\right)$
and the ”$k$-th step” partition function is defined by:
$Z_{N}(x,\beta_{N})=\mathbf{E}_{x}[e^{\sum_{n=t_{k-1}+1}^{t_{k}}\beta_{N}\left(w(n,S_{n})-\frac{1}{2}\beta^{2}\right)}],\text{ where $t_{k}=N^{\frac{k}{M}}$}$
Theorem 2.1 in [9] suggests that the limit (when $N\to\infty$) of increments $X_{k}$ are centered random variables with strictly increasing variance (in $k$). Heuristically, one can think of $M_{N}(x)$ as a branching random walk where the increments are given by the $X_{k}(x)$’s. Indeed, we can see that these increments are independent in $k$, satisfy that $X_{k}(x)$ and $X_{k}(y)$ are correlated (having high covariance) for small |xy||x-y| ($x,y\in\mathbb{Z}^{2}$), and decorrelate (having low variance) as |xy||x-y| grows large. In this analogy, the root of the tree corresponds to level $k=M$, and the leaves correspond to level $k=1$. Consequently, the asymptotic behavior of $h_{N}(x)$ should be governed by a (Gaussian) branching random walk model with decreasing variance.
It would be interesting to establish a rigorous connection between the time-inhomogeneous BRWs with decreasing variance studied in this paper and the log-partition function of two-dimensional directed polymers. In fact, this heuristic has been confirmed in [10] for the extreme values of $h_{N}(x)$, and raises an intriguing open problem: whether the two-dimensional polymer measure defined by:
$\mu_{N}^{\beta}:=\frac{e^{\beta h_{N}}}{\mathbb{E}\left[e^{\beta h_{N}}\right]}$
converges to a GMC measure. The universality of partition functions in the present paper (Theorem 3.3) suggests an affirmative answer to this problem (posed in [10]), though additional technical work is required to prove this rigorously.
4.3 Other Related Models
In this final section, we would other related models with time-inhomogeneous BRWs discussed in the present paper..
An interesting implication of Theorem 3.3 is that if we embed the (binary) BRW into the unit interval $[0,1]$, the covariance structure $K(x,y)$ (for $x,y\in[0,1]$) differs from that of GMC (which takes the form $\log|x-y|+O(1)$): it is not log-correlated. The universality behavior we established enables us to investigate whether we can construct a GMC-type object beyond log-correlated fields, which defines a variant of GMC measure. In this setting, the covariance kernel takes form of $F(x,y)+O(1)$, where $F$ is a positive, strictly non-increasing function satisfying $F(x,y)<\log|x-y|$ for all pairs of $x,y$.
Additionally, the results regarding to time-inhomogeneity of the increments discussed in the present paper raises a more fundamental questions: does addition of environmental fluctuations impact phase transitions (or other qualitative behavior) in disordered systems?
A particularly natural extension is the continuous time analogy of BRWs: the time-inhomogeneous Branching Brownian Motions (BBM). BBM were studied earlier than BRWs and are of independent interest (see [15] for a precise introduction for its time-inhomogeneous counterpart). Indeed, many results in branching random walks are known in the case of the branching Brownian motion, and the modifications of methods are more or less painless (and sometimes even neater). Conversely, once a result is established in Branching Brownian motions, it was often believed that a corresponding result holds in the context of BRWs (though not always straightforward to extend).
On the other hand, the model of BRWs has recently been viewed as a specific example of directed polymers on infinite graphs (see [11] for an introduction to this topic), and we hope our results can be extended to other related models, such as $\lambda$-biased BRWs on Galton-Watson trees. Finally, it is also interesting to extend the results to environments that are not independent: i.e., time-correlated random fields. There are more interesting dependence of time, compared to monotonic and macroscopic decay in variance discussed in the present paper in the modern literature. For example, the study of the influence of environmental fluctuations, or stochastic impurities, on directed polymers has been initiated in [7]. The authors sincerely hope that more interesting properties and phenomena will be discovered in the time-inhomogeneous settings of these models.